# Xing Huang a), Feng-Yu Wang a)

### v2

@@ Abstract
The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies a standard integrability condition in time-space, and may be discontinuous in the distance induced by any Dini function. The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the quadratic Wasserstein distance.
@@ 1 Introduction
$b:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
are measurable. Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 22] and the survey [11].
@@
$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|^{2}+\mathbb{W}_{2}(\mu,\nu)^{2}),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
@@
If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$ (see Definition 3.1 below), the following Bismut type formula has been established in [29] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
@@
rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable. This result encourages us to establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with coefficients singular in all time-spatial-distribution variables.
@@
where $\alpha$ is the square root of a Dini function, i.e. $\alpha$ is concave with $\alpha(0)=0,\alpha^{\prime}(s)>0$ for $s>0$, and $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s<\infty$ (Dini condition for $\alpha^{2}$). Thus, the continuity in the distribution variable is even weaker than the Dini continuity, so that the existing study in the literature is considerably improved.
To make this improvement, we develop a new technique by establishing a-priori derivative estimates and Bismut formula for the decoupled singular SDEs. The key point is to find a constant $c>0$ such that
(1.2) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
and for the drift of type $b_{t}(x,\mu(V))$ with $[V]_{\alpha}:=\sup_{x\neq y}\frac{|V(x)-V(y)|}{\alpha(|x-y|)}\leq 1$,
(1.3) $\|D^{I}P_{t}V(\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},$
where $D^{I}$ is the intrinsic derivative (see Definition 3.1 below).
@@ 2 Log-Harnack Inequality
Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for distributions in $\mathscr{P}_{2}$, but the drift may be discontinuous in $\mathbb{W}_{k}$ for any $k>0$. We first state the concrete assumption and the main result on the log-Harnack inequality, then present a complete proof in a separate subsection.
@@ 2.1 Assumption and main result
Let $\alpha$ be in the following class
$\mathscr{A}:=\bigg{\{}\alpha:[0,\infty)\rightarrow[0,\infty)\ \text{is\ increasing\ and\ concave,\ }\alpha(0)=0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}},$
where $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$. For a (real or Banach valued) function $f$, let
$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}$
be its continuity modulus in $\alpha$. Define the Wasserstein distance induced by $\alpha$:
$\mathbb{W}_{\alpha}(\mu,\nu):=\sup_{[f]_{\alpha}\leq 1}|\mu(f)-\nu(f)|,\ \ \mu,\nu\in\mathscr{P}_{\alpha}:=\big{\{}\mu\in\mathscr{P}:\mu(\alpha(|\cdot|))<\infty\big{\}},$
where $f$ are real functions and $\mu(f)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\mu$.
By the concavity of $\alpha$, $\mathbb{W}_{\alpha}$ is a complete distance on $\mathscr{P}_{\alpha}$, $\mathscr{P}_{k}\subset\mathscr{P}_{\alpha}$ for $k\geq 1$, and
@@
Since $\alpha$ is increasing with $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s>0,$ we have $\alpha(1)>0$ so that the second estimate in (2.1) with $r=t^{-1}$ yields
@@
$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2},\ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
@@
    There exist $\alpha\in\mathscr{A}$, $k\in(1,\infty),K\in(0,\infty),l\in\mathbb{N},$ and
@@ 3. $(A_{2})$
We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
@@
So, by [28, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
@@
(2.4) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{n}^{n}\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{n}\Big{]}\leq c_{n}(1+\|\gamma\|_{n}^{n}),\ \ n\geq 1.$
@@
Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfy
@@
(2.8) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+\nabla_{b_{t}(\cdot,P_{t}^{*}\gamma)}.$
By [32, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.7) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
@@
Moreover, for any $\gamma\in\mathscr{P}_{k}$, consider
(2.11) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
@@
Let $\sigma$ and $b$ satisfy (A). The the following assertions hold.
(1) The SDE (2.6) is well-posed and for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
Moreover, letting $\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}$ and $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
then the Bismut formula for $P^{\gamma}_{t}$ holds:
$\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\ \ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].$
(2) For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(2.15) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
@@
By [29, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (A) implies (1). It remains to prove (2.15). To this end, we use Zvonkin’s transform.
By Itô’s formula (see [32, Lemma 3.3]), we derive
@@
$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t})-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
@@
This together with (2.13), (2.16) and Gronwall’s inequality implies (2.15). ∎
@@
(2.21) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla_{b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
@@
Combining this with (2.20), (2.21), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [30, Lemma 2.1 and Lemma 4.1]), we derive
@@
by the stochastic Gronwall inequality (see [31, Lemma 3.7]), we find a constant $c_{2}>1$ such that
@@
To build the estimate (1.2), we present the following Hölder inequality for concave functions.
Lemma 2.5.
Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. For any non-negative random variables $\xi$ and $\eta$,
$\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
Consequently, for any random variable $\xi$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\eta$ with $\mathbb{E}[\eta]=0$,
(2.22) $\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi-y\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,y\in\mathbb{R}^{d}.$
Proof.
Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$. It suffices to prove for $\mathbb{E}[|\eta|^{p}]\in(0,\infty)$. Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using (2.1), we obtain
$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[|\eta|]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[|\eta|]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[|\eta|]\alpha\bigg{(}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
$\leq\mathbb{E}[|\eta|]\bigg{\{}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\ \alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[|\eta|^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
Then the second inequality follows by noting that $\mathbb{E}[\eta]=0$ implies
$\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\xi)-f(y)\}\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\xi-y|)|\eta|].$
@@
The following lemma provides an extension to (1.2), which combined with Lemma 2.4 implies the desired (2.17).
@@
$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
@@
It suffices to prove the first estimate, since it implies (2.24) according to Lemma 2.4.

### v4

@@ Abstract
The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies an integrability condition in time-space, and the continuity in distribution may be weaker than Dini. The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the 2-Wasserstein distance.
@@ 1 Introduction
$b:[0,T]\times\mathbb{R}^{d}\times\tilde{\mathscr{P}}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
are measurable for some non-empty subspace $\tilde{\mathscr{P}}\subset\mathscr{P}$ equipped with a complete distance $\tilde{\rho}$. Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 23] and the survey [11].
@@
$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|+\mathbb{W}_{2}(\mu,\nu)),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
@@
If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$, the following Bismut type formula has been established in [30] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
@@
for some $\varepsilon\in(0,1)$ rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable. This result encourages us to establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with coefficients singular in all time-spatial-distribution variables.
@@
where $\alpha$ is the square root of a Dini function, i.e. it belongs to class
[…]
$\alpha(0)=0,\ \alpha(r)>0\ \text{for}\ r>0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}}.$
Noting that $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$, the continuity in the distribution variable is even weaker than Dini, so that the existing study in the literature is considerably improved.
The log-Harnack inequality is established in Section 2, where a key step is to derive the estimate (Lemma 2.6 for $k=2$):
$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},t\in(0,T]$
for some constant $c>0$.
The Bismut formula for the intrinsic derivative of $P_{t}f$ is presented in Section 3, for which we develop new techniques to control the intrinsic derivative $D^{I}$ and the extrinsic derivative $D^{E}$ of the drift term in the distribution variable (Theorem 3.3(1)):
$\|D^{I}P_{t}[D^{E}b_{t}(y,\nu)(\cdot)](\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
@@ 2 Log-Harnack Inequality
Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{2},\mathbb{W}_{2})$, but the drift may be not Lipschitz continuous in $\mathbb{W}_{k}$ for any $k>0$. We first state the concrete assumption and the main result on the log-Harnack inequality, then present a complete proof in a separate subsection.
@@ 2.1 Assumption and main result
Let $\alpha\in\mathscr{A}$. Then it holds
@@
The second estimate in (2.1) with $r=t^{-1}$ yields
@@
$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2}:\ \ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
@@
    Let $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{k},\mathbb{W}_{k})$ for some $k\in(1,\infty)$. There exist $K\in(0,\infty),l\in\mathbb{N},\alpha\in\mathscr{A}$ and
@@ 3. $(A_{2})$
We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Let $[\cdot]_{\alpha}$ be the $\alpha$-continuity modulus defined by
$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}.$
Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq\alpha(1)(1+|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
@@
So, by [29, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
(2.4) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
@@
(2.5) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{k}^{k}=\sup_{t\in[0,T]}(P_{t}^{*}\gamma)(|\cdot|^{k})\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{k}\Big{]}\leq c_{k}(1+\|\gamma\|_{k}^{k}).$
@@
Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfy
@@
By [29, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, this SDE is well-posed and (2.4) also holds for $X_{t}^{x,\gamma}$ in place of $X_{t}$, i.e. for any $n\geq 1$ there exists a constant $c_{n}(\gamma)>0$ such that
(2.8) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x,\gamma}|^{n}]\leq c_{n}(\gamma)(1+|x|^{n}),\ \ x\in\mathbb{R}^{d}.$
Let $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
We first present the following generalized Hölder inequality with a concave function $\alpha$.
Lemma 2.3.
Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. Then for any non-negative random variables $\xi$ and $\eta$,
(2.9) $\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
Consequently, for any random variable $\bar{\xi}$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\bar{\eta}$ with $\mathbb{E}[\bar{\eta}]=0$,
(2.10) $\big{\|}\mathbb{E}[f(\bar{\xi})\bar{\eta}]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\bar{\eta}\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\bar{\xi}-z\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,z\in\mathbb{R}^{d}.$
Proof.
Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$. It suffices to prove for $\mathbb{E}[\eta^{p}]\in(0,\infty)$. Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using the second inequality in (2.1), we obtain
$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[\eta]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[\eta]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[\eta]\alpha\bigg{(}\frac{(\mathbb{E}[\eta^{p}])^{\frac{1}{p}}}{\mathbb{E}[\eta]}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
$\leq\mathbb{E}[\eta]\bigg{\{}\frac{(\mathbb{E}[\eta^{p}])^{\frac{1}{p}}}{\mathbb{E}[\eta]}\ \alpha\Big{(}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[\eta^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
Then the second inequality follows by noting that $\mathbb{E}[\bar{\eta}]=0$ implies
$\big{\|}\mathbb{E}[f(\bar{\xi})\bar{\eta}]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\bar{\xi})-f(z)\}\bar{\eta}]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\bar{\xi}-z|)|\bar{\eta}|].$
Therefore, the proof is completed. ∎
@@
(2.12) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+b_{t}(\cdot,P_{t}^{*}\gamma)\cdot\nabla.$
By [33, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.11) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
@@
Moreover, for any $\gamma\in\mathscr{P}_{k}$, $t\in[0,T]$, consider
(2.15) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ \theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
@@
Let $\sigma$ and $b$ satisfy (A). Then the following assertions hold.
  1. $(1)$
    For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
    (2.18) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k}.$
  2. $(2)$
    For any $\alpha\in\mathscr{A}$, there exists a constant $c>0$ such that the gradient estimate holds:
    (2.19) $\begin{split}&|\nabla P^{\gamma}_{t}f|(x):=\limsup_{|y-x|\rightarrow 0}\frac{|P^{\gamma}_{t}f(y)-P^{\gamma}_{t}f(x)|}{|y-x|}\\ &\leq\frac{c\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \ [f]_{\alpha}\leq 1,\ x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
@@ 2. $(2)$
(1) We will use Zvonkin’s transform defined in (2.14). By Itô’s formula (see [33, Lemma 3.3]), we derive
@@
$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x),P_{t}^{*}\gamma)\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
@@
This together with (2.17), (2.20) and Gronwall’s inequality implies (2.18).
(2) For any measurable $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ with $[f]_{\alpha}\leq 1$, take
$f_{n}:=[(-n)\lor f]\land n,\ \ n\geq 1.$
By an approximation technique, it is sufficient to prove (2.19) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ with $[f]_{\alpha}\leq 1$. According to [33, Theorem 4.1], there exists a constant $c_{0}>0$ such that for any $\gamma\in\mathscr{P}_{k}$, the log-Harnack inequality
$P_{t}^{\gamma}\log f(x)\leq\log P_{t}^{\gamma}f(y)+\frac{c_{0}}{t}|x-y|^{2},\ \ x,y\in\mathbb{R}^{d},t\in(0,T],f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d})$
holds, so that [27, Proposition 1.3.8] implies
$|\nabla P^{\gamma}_{t}f|\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{P_{t}^{\gamma}|f|^{2}\}^{\frac{1}{2}},\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),t\in(0,T],\gamma\in\mathscr{P}_{k}.$
Observe that for any $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ with $[f]_{\alpha}\leq 1$,
$|\nabla P^{\gamma}_{t}f|(x)$ $\leq\inf_{z\in\mathbb{R}}\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{P_{t}^{\gamma}(|f-z|^{2})(x)\}^{\frac{1}{2}}$
(2.21) $\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{\mathbb{E}(\alpha(|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|)^{2})\}^{\frac{1}{2}}$
$\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\alpha(\{\mathbb{E}(|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{2})\}^{\frac{1}{2}}),\ \ x\in\mathbb{R}^{d},t\in(0,T],$
where in the last step, we used (2.9) for $\eta=\alpha(\xi)$ with $\xi=|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|$ and $p=2$. Therefore, (2.19) follows from (2.2), (2.18) and (2.1). ∎
@@
(2.26) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})[b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)]\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
@@
Combining this with (2.25), (2.26), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [31, Lemma 2.1 and Lemma 4.1]), we derive
@@
by the stochastic Gronwall inequality (see [32, Lemma 3.7]), we find a constant $c_{2}>1$ such that
@@
The following lemma provides a regularity estimate on $P_{t}^{*}$, which together with Lemma 2.5 implies the desired (2.22).
@@
(2.28) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
@@
By Lemma 2.5, (2.29) follows from (2.28) and the fact $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}\text{\rm{d}}t<\infty$. So, we only need to prove (2.28).